OpenStudy (anonymous):

f(-7) if f(x) = 2x + 8.

OpenStudy (igreen):

Just plug in -7 for x: f(-7) = 2(-7) + 8 f(-7) = -14 + 8 f(-7) = -6

OpenStudy (anonymous):

-6. can someone second?

OpenStudy (igreen):

Welcome to Open Study! You can give medals by clicking 'Best Response'. @paige56476